# Structural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengths

### v1

@@
Structural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengths
Michele Pizzardo Structural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengthsStructural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengthsStructural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengthsStructural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengths mpizzardo@asiaa.sinica.edu.tw    Ivana Damjanov Structural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengthsStructural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengths    Jubee Sohn Structural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengthsStructural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengths    Margaret J. Geller Structural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengthsStructural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengths
@@ 1 Introduction
The observed evolution in galaxy morphological properties with redshift is a fundamental test of theories that describe galaxy mass assembly. For galaxies that have just halted global star formation, sizes and central concentrations derived from light profiles at different rest-frame wavelengths star constrain physical mechanisms that drive the evolution. For galaxies with old stellar populations, structural parameters estimated at multiple fixed rest-frame wavelengths constrain the channels for growth in the quiescent phase.
Size growth at fixed stellar mass is the most widely studied aspect of the structural evolution of quiescent galaxies. Evaluation of the size growth of the quiescent population over a broad redshift range is based on the size-stellar mass relation, a projection of the stellar mass fundamental plane (e.g., Hyde & Bernardi 2009; Bezanson et al. 2013; Zahid et al. 2016). For quiescent galaxies with stellar masses $\gtrsim 2\times 10^{10}$, this relation has the form $R_{e}=R_{0}(M_{\star}/M_{0})^{\alpha}$, where $R_{e}$ is the radius enclosing half of the total light profile of a galaxy (i.e., half-light radius), $M_{\star}$ is galaxy stellar mass, and $R_{0}$ the characteristic half-light radius (a proxy for size) of galaxies with $M_{\star}=M_{0}$. Based mostly on galaxy light profiles in a single photometric band, previous investigations suggest that the slope $\alpha$ varies very little with redshift; in contrast, the normalization $R_{0}$ increases substantially. From a redshift of $\sim 2$ to the present, systems with masses $\gtrsim 2\times 10^{10}$ grow of a factor of 2.5-4 (e.g., Daddi et al. 2005; Trujillo et al. 2007; van Dokkum et al. 2010; Damjanov et al. 2011; Newman et al. 2012; Huertas-Company et al. 2013; van der Wel et al. 2014; Huertas-Company et al. 2015; Faisst et al. 2017; Damjanov et al. 2019; Mowla et al. 2019; Kawinwanichakij et al. 2021; Barone et al. 2022; Hamadouche et al. 2022; Damjanov et al. 2023; George et al. 2024; Euclid Collaboration et al. 2025a). Recent results from the James Webb Space Telescope (JWST) extend constraints on this evolution to $z\sim 3-4$ (Cutler et al. 2024; Ito et al. 2024; Kawinwanichakij et al. 2025).
Evolution of the Sérsic index, a proxy for the central concentration of galaxy light, probes the connection between galaxy mass assembly and the distribution of its stellar content with cosmic time. Single photometric band investigations at redshift $\lesssim 2.5$ (Huertas-Company et al. 2013; Patel et al. 2013; Lang et al. 2014; Whitaker et al. 2015; Martorano et al. 2025) demonstrate that the median Sérsic index of quiescent galaxies evolves slowly or not at all with redshift. Very high-mass galaxies $(M_{\star}>10^{11}$ ${}_{\odot})$ display the largest redshift dependence, with an increase from $n\sim 3.5$ to $n\sim 4.5$ as redshift decreases from $z\sim 1.5$ to $z\sim 0.5$.
@@
Investigators tackle the wavelength dependence of morphological parameters in different ways. Often, analyses adopt a single rest-frame wavelength to explore the bulk of the galaxy stellar mass content, typically $\sim 5000$ (e.g., van der Wel et al. 2014; Mowla et al. 2019; Kawinwanichakij et al. 2021). Galaxy light profiles at this wavelength are dominated by the light from low-mass stars that trace the total stellar mass of a galaxy (e.g., Bruzual & Charlot 2003). Alternatively, multi-wavelength studies use mass-weighted structural parameters based on measurements from galaxy light profiles in a fixed band multiplied by the profile of the stellar mass–to–light ratio ($M/L$) in that band from (spatially resolved) spectral energy distribution fitting (e.g., Suess et al. 2019a, b; Clausen et al. 2025). Some multi-band fitting algorithms provide automated extraction of morphological parameters as functions of wavelength (Häußler et al. 2013; Nedkova et al. 2021).
@@
The change in size and shape of galaxies with rest-frame wavelength for galaxy samples with precisely determined redshifts can trace differences in the distribution of their stellar populations as a function of different properties (age, metallicity). Red/infrared light traces the mass content of galaxies; in contrast, ultraviolet light (UV, $\sim(2000-4000)$) probes the distribution of recently formed and/or metal-poor stars (e.g., Bruzual & Charlot 2003). Wavelength-related variations in the structural properties (size, central concentration) of (globally) young quiescent galaxies directly reflect the dominant physical mechanisms that halt star formation in these systems (e.g., inside-out vs. inside-in quenching, Avila-Reese et al. 2018; Lin et al. 2019; Papaderos, Polychronis et al. 2022). Changes in structural parameters with wavelength for galaxies dominated by the oldest stellar populations track galaxy evolution during the quiescent phase (e.g., the growth in size via minor mergers, Naab et al. 2009; Nipoti et al. 2009; Hilz et al. 2012; Oser et al. 2012; Nipoti et al. 2012).
@@
The rest-frame wavelength dependence of galaxy morphological parameters have been studied in detail with large spectroscopic samples of quiescent systems at $z\lesssim 0.2$ (e.g., GAMA, Driver et al. 2009; Kelvin et al. 2012; Kennedy et al. 2016). At higher redshifts, $\sim 0.5-2$, the trends in structural properties with rest-frame wavelength are available only for small spectro-photometric samples from HST (e.g., Guo et al. 2011; van der Wel et al. 2014; Suess et al. 2020) and JWST (e.g., Clausen et al. 2025; Martorano et al. 2025) or for statistically large purely photometric samples (e.g., from Subaru/HSC, George et al. 2024). A study of the first statistically large dataset from the Euclid satellite includes measurements of morphological parameters in the visible and near-infrared filters for a photometric sample of $\sim 100,000$ quiescent galaxies at $z\lesssim 1$ (over $<0.5\%$ of the planned total survey area, Euclid Collaboration et al. 2025a).
@@
These studies show that, for massive quiescent galaxies ($M_{\star}\gtrsim 10^{10}\,M_{\odot}$) covering a broad redshift range, the galaxy size decreases and the Sérsic index increases with increasing wavelength. The redder light profiles of quiescent galaxies reflect the older stellar population and the bulk of the galaxy stellar mass. They cover smaller galactocentric radial ranges and are more centrally concentrated than their bluer counterparts dominated by the light from young and/or metal-poor stellar populations.
@@
The investigation of structural parameters with a large, mass-complete sample of quiescent galaxies including both high-quality multi-band imaging and spectroscopy in the redshift range $0.60.2<z<$ 0.6 is a missing link between studies with statistically large samples at $z\sim 0$ and results based on much smaller sets of spectroscopically-confirmed quiescent systems at $z>0.5$. Here we provide the first measurements covering this redshift gap. We extend studies of wavelength-dependent structural parameters with statistical samples of quiescent galaxies up to $z\sim 0.6$.
@@
We select the quiescent sample from the HectoMAP spectroscopic survey (Geller & Hwang 2015; Sohn et al. 2021, 2023), a dense red-selected redshift survey covering 55 square degrees of the northern sky. HectoMAP includes a complete mass-limited sample of $\sim 40,000$ quiescent systems at $0.60.2<z<$ 0.6. In addition to precise redshift measurements, medium resolution spectroscopy in the observed $3500-9000$ interval enables measurements of the spectral index $D_{n}4000$, a ratio of the flux in the continuum around the (rest-frame) $4000$ break and a proxy for average galaxy stellar population age. The HectoMAP field is included in the Subaru Hyper Suprime-Cam (HSC) Subaru Strategic Program (SSP) multi-band ($grizy$) imaging (Aihara et al. 2022). Dense spectroscopy combined with multi-wavelength high-resolution imaging makes the HectoMAP galaxy sample a unique tracer of the evolution in galaxy morphological parameters over the redshift range $0.60.2<z<$ 0.6 as a function of rest-frame wavelength, stellar mass, and average stellar population age.
@@
We summarize the HectoMAP survey in Sect. 2.1 and review the complete sample of quiescent galaxies in Sect. 2.2. We describe the HSC-SSP imaging in Sect. 3.1, the Sérsic fitting procedure in Sect. 3.2, and the results of multi-band fitting to the HectoMAP galaxies in Sect. 3.3. We describe the resulting wavelength dependence of size and Sérsic index in Sect. 4.1, and we investigate the dependence of these (anti-)correlations on $D_{n}4000$ in Sect. 4.2. In Sections 5.1 and 5.2, we investigate how the size–mass and Sérsic index–mass relations evolve in two rest-frame regimes, mapping structural changes in different stellar populations of quiescent systems. We conclude in Sect. 6.
@@ 2 Spectroscopy from MMT/Hectospec
We explore the wavelength dependence of the structural parameters of quiescent galaxies at redshifts 0.2 to 0.6 based on single-Sérsic profile fitting. We combine medium-resolution MMT/Hectospec spectroscopy and high-resolution multi-band Subaru/HSC imaging to select and characterize the sample.
Sect. 2.1 briefly outlines HectoMAP, the spectroscopic galaxy survey that provides the parent sample of quiescent galaxies. Sect. 2.2 describes the mass-complete sample of HectoMAP galaxies we analyze. Table 1 lists the main characteristics of the HectoMAP parent sample along with the selection that underlies the galaxy sample we use.
@@ 2.1 The HectoMAP galaxy redshift survey
HectoMAP is a large dense redshift survey of quiescent galaxies covering redshifts $0.60.2<z<$ 0.6 carried out with the Hectospec 300-fiber wide-field instrument on the 6.5-meter MMT (Fabricant et al. 2005). HectoMAP includes spectra for 95,403 galaxies with $r<23$ covering 54.64 deg2 over a narrow strip of the sky at $200<$ R.A. (deg) $<250$ and $42.5<$ Decl. (deg) $<44.0$ (first and second rows in Table 1). The median redshift of HectoMAP is $z=0.345$ (Geller et al. 2011; Geller & Hwang 2015; Sohn et al. 2021, 2023).
Table 1: Selection of the HectoMAP Subsample
@@
b We remove galaxies with $0.420.4<z<$ 0.42 because of night sky line contamination (Damjanov et al. 2024).
c HSC-SSP coverage of the HectoMAP field is $\sim 80\%$ for the $i$- and $y$-bands, and $100\%$ for the $g$-, $r$-, and $z$-bands (Sect. 3.1).
The photometric calibration of HectoMAP survey is based on the Sloan Digital Sky Survey (SDSS) DR16 (Ahumada et al. 2020). HectoMAP incudes two samples: a bright survey with $r_{petro,0}<20.5$ and a red color selection $(g-r)_{model,0}>1$, and a fainter survey with $20.5<r_{petro,0}\leq 21.3$ and red selection $(g-r)_{model,0}>1$ and $(r-i)_{model,0}>0.5$. The entire survey contains 74,365 galaxies (third and fourth rows in Table 1). Here $r_{petro,0}$ refers to the SDSS Petrosian magnitude corrected for Galactic extinction. HectoMAP color selection is based on the extinction-corrected SDSS model magnitudes. For the faint portion, the $(r-i)_{model,0}$ cut removes low redshift objects. There is an additional surface brightness limit $r_{fiber,0}<22.0$ throughout the HectoMAP survey because red objects below this limit yield very low signal-to-noise MMT spectra. This limiting $r_{fiber,0}$ magnitude corresponds to the extinction-corrected flux within the Hectospec fiber aperture (with $0\aas@@fstack{\prime\prime}75$ radius).
Sohn et al. (2021, 2023) describe two HectoMAP data releases that include the entire redshift survey. The releases include the redshifts, spectral index $D_{n}4000$ measurements, and stellar masses along with the match to the SDSS photometry. We summarize these data.
@@
In Sohn et al. (2023) we compute stellar masses for 99.8% of HectoMAP galaxies based on SDSS DR16 $ugriz$ photometry. We use the Le Phare software package (Arnouts et al. 1999; Ilbert et al. 2006) that incorporates stellar population synthesis models of Bruzual & Charlot (2003), a Chabrier (2003) initial mass function, and a Calzetti et al. (2000) extinction law. We explore two metallicities (0.4 and 1 solar) and a set of exponentially decreasing star formation rates. We derive the mass-to-light ratio from the best model and use this ratio to convert the observed luminosity into stellar mass.
@@
The HectoMAP data releases also include the spectral indicator $D_{n}{4000}$, a ratio of fluxes around the $4000$ break. $D_{n}{4000}$ is a proxy for the average age of the galaxy stellar population (Kauffmann et al. 2003; Damjanov et al. 2024). Following the definition from Balogh et al. (1999), we compute the ratio of fluxes in two spectral windows: (3850–3950) Å and (4000–4100) Å. The median S/N at (3850–4100) Å is $\sim 4.5$. We measure $D_{n}4000$ for $\sim 96.5\%$ of the color-selected HectoMAP galaxies with mass measurement (sixth row of Table 1).
Based on the estimated range of spectral indices from a sample of synthetic galaxy spectra (Kauffmann et al. 2003) and the signal-to-noise of HectoMAP spectra (Table 1 and Figure 7 of Damjanov et al. 2023), we evaluate the reliability of the $D_{n}{4000}$ measurements for the parent galaxy sample. The criteria we use are: (1) $D_{n}4000>0$, (2) $D_{n}4000<3$, and (3) $\Delta D_{n}4000<0.2D_{n}4000$.
For galaxies with $D_{n}{4000}<2.2$, the measurement error $\Delta D_{n}4000$ does not depend on $D_{n}{4000}$; in contrast, $\Delta D_{n}4000$ increases with $D_{n}{4000}$ for galaxies with $D_{n}{4000}>2.2$. We resample $D_{n}{4000}$ for galaxies with $32.2<D_{n}{4000}<$ 3 by first computing the average ${\Delta D_{n}4000}$$\overline{\Delta D_{n}4000}$ – for all of the galaxies in this subset. Then for each galaxy $i$ from the subset we randomly generate a set of corrections $\varepsilon_{i_{j}}$, where $j\in[1,100]$, from a normal distribution with zero mean and dispersion equal to $\overline{\Delta D_{n}4000}$. Using the average $\bar{\varepsilon_{i}}$ of these corrections, we compute ${D_{n}4000}_{i}^{new}={D_{n}4000}_{i}+\bar{\varepsilon_{i}}$. We retain galaxy $i$ with its corresponding ${D_{n}4000}_{i}^{new}$ in the sample if this new value is smaller than 2.2. This correction procedure for $D_{n}{4000}$ applies to only $\sim 4.4\%$ of the galaxies with $D_{n}4000$ measurements; we remove $\sim 50\%$ of galaxies with initial index values $32.2<D_{n}4000<$ 3. In Sect. 2.2 we use $D_{n}{4000}$ to select quiescent HectoMAP galaxies.

### v2

# Structural evolution of quiescent galaxies at intermediate redshifts at UV and red rest-frame wavelengths

@@
Structural evolution of quiescent galaxies at intermediate redshifts at UV and red rest-frame wavelengths
Michele Pizzardo    Ivana Damjanov    Jubee Sohn    Margaret J. Geller
@@ 1 Introduction
The observed evolution in galaxy morphological properties with redshift is a fundamental test of theories that describe galaxy mass assembly. For galaxies that have just halted their global star formation processes, the sizes and central concentrations derived from light profiles at different rest-frame wavelengths can constrain physical mechanisms driving their evolution. For galaxies with old stellar populations, estimates of the structural parameters at multiple fixed rest-frame wavelengths can constrain the channels for growth in the quiescent phase.
Size growth at fixed stellar mass is the most widely studied aspect of the structural evolution of quiescent galaxies. The evaluation of the size growth of the quiescent population over a broad redshift range is based on the size-stellar mass relation, a projection of the stellar mass fundamental plane (e.g., Hyde & Bernardi 2009; Bezanson et al. 2013; Zahid et al. 2016). For quiescent galaxies with stellar masses $\gtrsim 2\times 10^{10}$M, this relation takes the form of $R_{e}=R_{0}(M_{\star}/M_{0})^{\alpha}$, where $R_{e}$ is the radius enclosing half of the total light profile of a galaxy (i.e., half-light radius), $M_{\star}$ is the galaxy stellar mass, and $R_{0}$ is the characteristic half-light radius (a proxy for size) of galaxies with $M_{\star}=M_{0}$. Based primarily on galaxy light profiles in a single photometric band, previous investigations suggest that the slope, $\alpha$, varies very little with redshift; in contrast, the normalization, $R_{0}$, increases substantially. From a redshift of $\sim 2$ to the present, systems with masses $\gtrsim 2\times 10^{10}$M grow by a factor of 2.5-4 (e.g., Daddi et al. 2005; Trujillo et al. 2007; van Dokkum et al. 2010; Damjanov et al. 2011; Newman et al. 2012; Huertas-Company et al. 2013; van der Wel et al. 2014; Huertas-Company et al. 2015; Faisst et al. 2017; Damjanov et al. 2019; Mowla et al. 2019; Kawinwanichakij et al. 2021; Barone et al. 2022; Hamadouche et al. 2022; Damjanov et al. 2023; George et al. 2024; Euclid Collaboration et al. 2025b). Recent results from the James Webb Space Telescope (JWST) extend the constraints on this evolution to higher redshift (Cutler et al. 2024; Ito et al. 2024; Martorano et al. 2024; Hamadouche et al. 2025; Kawinwanichakij et al. 2025).
[…]
@@
Investigators tackle the wavelength dependence of morphological parameters in different ways. Often, analyses adopt a single rest-frame wavelength to explore the bulk of the galaxy stellar mass content, typically $\sim 5000$ Å (e.g., van der Wel et al. 2014; Mowla et al. 2019; Kawinwanichakij et al. 2021). Galaxy light profiles at this wavelength are dominated by the light from low-mass stars that trace the total stellar mass of a galaxy (e.g., Bruzual & Charlot 2003). Alternatively, multiwavelength studies use mass-weighted structural parameters based on measurements from galaxy light profiles in a fixed band multiplied by the profile of the stellar-mass-to-light ratio ($M/L$) in that band from (spatially resolved) spectral energy distribution fittings (e.g., Suess et al. 2019a, b; Clausen et al. 2025). Some multiband fitting algorithms provide automated extraction of morphological parameters as functions of wavelength (Häußler et al. 2013; Nedkova et al. 2021).
@@
The change in the size and shape of galaxies with rest-frame wavelengths for galaxy samples with precisely determined redshifts can trace differences in the distribution of their stellar populations as a function of different properties (age, metallicity). Red, infrared light traces the mass content of galaxies; in contrast, ultraviolet light (UV, $\sim(2000-4000)$ Å) probes the distribution of recently formed and/or metal-poor stars (e.g., Bruzual & Charlot 2003). Wavelength-related variations in the structural properties (size, central concentration) of (globally) young quiescent galaxies directly reflect the dominant physical mechanisms that halt star formation in these systems (e.g., inside-out vs. inside-in quenching, Avila-Reese et al. 2018; Lin et al. 2019; Papaderos, Polychronis et al. 2022). Changes in structural parameters with wavelength for galaxies dominated by the oldest stellar populations track galaxy evolution during the quiescent phase (e.g., the growth in size via minor mergers, Naab et al. 2009; Nipoti et al. 2009; Hilz et al. 2012; Oser et al. 2012; Nipoti et al. 2012).
@@
The rest-frame wavelength dependence of galaxy morphological parameters has been studied in detail with large spectroscopic samples of quiescent systems at $z\lesssim 0.2$ (e.g., GAMA, Driver et al. 2009; Kelvin et al. 2012; Kennedy et al. 2016). At higher redshifts ($\sim 0.5-2$) the trends in the structural properties with rest-frame wavelength are available only for small spectro-photometric samples from HST (e.g., Guo et al. 2011; van der Wel et al. 2014; Suess et al. 2020) and JWST (e.g., Clausen et al. 2025) or for larger purely photometric samples (e.g., from Subaru/HSC or JWST, Cutler et al. 2024; George et al. 2024; Martorano et al. 2024, 2025). A study of the first statistically large dataset from the Euclid satellite includes measurements of morphological parameters in the visible and near-infrared filters for a photometric sample of $\sim 100,000$ quiescent galaxies at $z\lesssim 1$ (over $0.5\%$ of the planned total survey area, Euclid Collaboration et al. 2025b).
@@
These studies show that for massive quiescent galaxies ($M_{\star}\gtrsim 10^{10}\,M_{\odot}$) covering a broad redshift range, the galaxy size decreases and the Sérsic index increases with increasing wavelength. The redder light profiles of quiescent galaxies reflect the older stellar population and the bulk of the galaxy stellar mass. They cover smaller galactocentric radial ranges and are more centrally concentrated than their bluer counterparts dominated by the light from young and/or metal-poor stellar populations.
@@
The investigation of structural parameters with a large, mass-complete sample of quiescent galaxies including both high-quality multiband imaging and spectroscopy in the redshift range $0.2<z<0.6$ is a missing link between studies with statistically large samples at $z\sim 0$ and results based on much smaller sets of spectroscopically confirmed quiescent systems at $z>0.5$. Here, we provide the first measurements covering this redshift gap. We extend studies of wavelength-dependent structural parameters with statistical samples of quiescent galaxies up to $z\sim 0.6$.
@@
We selected the quiescent sample from the HectoMAP spectroscopic survey (Geller & Hwang 2015; Sohn et al. 2021, 2023), a dense red-selected redshift survey covering 55 square degrees of the northern sky. HectoMAP includes a complete mass-limited sample of $\sim 40,000$ quiescent systems at $0.2<z<0.6$. In addition to precise redshift measurements, medium-resolution spectroscopy in the observed 350090003500-9000 Å interval enables measurements of the spectral index $D_{n}4000$, which is a ratio of the flux in the continuum around the (rest-frame) $4000$ Å break and a proxy for average galaxy stellar population age. The HectoMAP field is included in the Subaru Hyper Suprime-Cam (HSC) Subaru Strategic Program (SSP) multiband (grizygrizy) imaging (Aihara et al. 2022). Dense spectroscopy combined with multiwavelength high-resolution imaging makes the HectoMAP galaxy sample a unique tracer of the evolution in galaxy morphological parameters over the redshift range of $0.2<z<0.6$ as a function of rest-frame wavelength, stellar mass, and average stellar population age.
@@
We summarize the HectoMAP survey in Sect. 2.1 and review the complete sample of quiescent galaxies in Sect. 2.2. We describe the HSC-SSP imaging in Sect. 3.1, the Sérsic fitting procedure in Sect. 3.2, and the results of multiband fitting to the HectoMAP galaxies in Sect. 3.3. We describe the resulting wavelength dependence of size and Sérsic index in Sect. 4.1, and we investigate the dependence of these (anti-)correlations on $D_{n}4000$ in Sect. 4.2. In Sects. 5.1 and 5.2 we investigate how the size–mass and Sérsic index–mass relations evolve in two rest-frame regimes, mapping structural changes in different stellar populations of quiescent systems. We present our conclusions in Sect. 6.
@@ 2 Spectroscopy from MMT/Hectospec
We explored the wavelength dependence of the structural parameters of quiescent galaxies at redshifts 0.2 to 0.6 based on single-Sérsic profile fitting. We combined medium-resolution MMT/Hectospec spectroscopy and high-resolution multiband Subaru/HSC imaging to select and characterize the sample.
Section 2.1 briefly outlines HectoMAP, the spectroscopic galaxy survey that provides the parent sample of quiescent galaxies. Section 2.2 describes the mass-complete sample of HectoMAP galaxies we analyze. Table 1 lists the main characteristics of the HectoMAP parent sample along with the selection that underlies the galaxy sample we use.
@@ 2.1 The HectoMAP galaxy redshift survey
HectoMAP is a large dense redshift survey of quiescent galaxies covering redshifts $0.2<z<0.6$ carried out with the Hectospec 300-fiber wide-field instrument on the 6.5-meter MMT (Fabricant et al. 2005). HectoMAP includes spectra for 95,403 galaxies with $r<23$ covering 54.64 deg2 over a narrow strip of the sky at $200<$ R.A. (deg) <250<250 and $42.5<$ Decl. (deg) <44.0<44.0 (first and second rows in Table 1). The median redshift of HectoMAP is $z=0.345$ (Geller et al. 2011; Geller & Hwang 2015; Sohn et al. 2021, 2023).
Table 1: Selection of the HectoMAP Subsample.
@@
b We remove galaxies with $0.4<z<0.42$ because of night sky line contamination (Damjanov et al. 2025).
c HSC-SSP coverage of the HectoMAP field is $\sim 80\%$ for the $i$ and $y$ bands, and $100\%$ for the $g$, $r$, and $z$ bands (Sect. 3.1).
The photometric calibration of HectoMAP survey is based on the Sloan Digital Sky Survey (SDSS) DR16 (Ahumada et al. 2020). HectoMAP incudes two samples: a bright survey with $r_{petro,0}<20.5$ and a red color selection $(g-r)_{model,0}>1$, and a fainter survey with $20.5<r_{petro,0}\leq 21.3$ and red selection $(g-r)_{model,0}>1$ and $(r-i)_{model,0}>0.5$. The entire survey contains 74,365 galaxies (third and fourth rows in Table 1). Here $r_{petro,0}$ refers to the SDSS Petrosian magnitude corrected for Galactic extinction. HectoMAP color selection is based on the extinction-corrected SDSS model magnitudes. For the faint portion, the $(r-i)_{model,0}$ cut removes low redshift objects. There is an additional surface brightness limit $r_{fiber,0}<22.0$ throughout the HectoMAP survey because red objects below this limit yield very low signal-to-noise ratio MMT spectra. This limiting $r_{fiber,0}$ magnitude corresponds to the extinction-corrected flux within the Hectospec fiber aperture (with $0\aas@@fstack{\prime\prime}75$ radius).
Sohn et al. (2021, 2023) described two HectoMAP data releases that include the entire redshift survey. These releases include the redshifts, spectral index $D_{n}4000$ measurements, and stellar masses along with the match to the SDSS photometry. We summarize these data below.
@@
In Sohn et al. (2023), we computed the stellar masses for 99.8% of HectoMAP galaxies based on SDSS DR16 ugrizugriz photometry. We used the Le Phare software package (Arnouts et al. 1999; Ilbert et al. 2006) that incorporates stellar population synthesis models of Bruzual & Charlot (2003), a Chabrier (2003) initial mass function, and a Calzetti et al. (2000) extinction law. Following the approach from our earlier studies of galaxies at $z<1$ (e.g., Geller et al. 2014), we explored two metallicities (0.4 and 1 solar) and a set of exponentially decreasing star formation rates. We derived the mass-to-light ratio from the best model and used this ratio to convert the observed luminosity into stellar mass.
@@
The HectoMAP data releases also include the spectral indicator $D_{n}{4000}$, a ratio of fluxes around the $4000$ Å break. $D_{n}{4000}$ is a proxy for the average age of the galaxy stellar population (Kauffmann et al. 2003; Damjanov et al. 2025). Following the definition from Balogh et al. (1999), we compute the ratio of fluxes in two spectral windows: (3850–3950) Å and (4000–4100) Å. The median S/N at (3850–4100) Å is $\sim 4.5$. We measure $D_{n}4000$ for $\sim 96.5\%$ of the color-selected HectoMAP galaxies with mass measurement (sixth row of Table 1).
Based on the estimated range of spectral indices from a sample of synthetic galaxy spectra (Kauffmann et al. 2003) and the signal-to-noise of HectoMAP spectra (Table 1 and Fig. 7 of Damjanov et al. 2023), we evaluate the reliability of the $D_{n}{4000}$ measurements for the parent galaxy sample. The criteria we use are: (1) $D_{n}4000>0$, (2) $D_{n}4000<3$, and (3) $\Delta D_{n}4000<0.2D_{n}4000$.
For galaxies with $D_{n}{4000}<2.2$, the measurement error $\Delta D_{n}4000$ does not depend on $D_{n}{4000}$; in contrast, $\Delta D_{n}4000$ increases with $D_{n}{4000}$ for galaxies with $D_{n}{4000}>2.2$. We resample $D_{n}{4000}$ for galaxies with $2.2<D_{n}{4000}<3$ by first computing the average $\Delta D_{n}4000$$\overline{\Delta D_{n}4000}$ – for all of the galaxies in this subset. Then for each galaxy, $i$, from the subset we randomly generate a set of corrections $\varepsilon_{i_{j}}$, where $j\in[1,100]$, from a normal distribution with zero mean and dispersion equal to $\overline{\Delta D_{n}4000}$. Using the average $\bar{\varepsilon_{i}}$ of these corrections, we computed ${D_{n}4000}_{i}^{new}={D_{n}4000}_{i}+\bar{\varepsilon_{i}}$. We retained galaxy $i$ with its corresponding ${D_{n}4000}_{i}^{new}$ in the sample if this new value is smaller than 2.2. This correction procedure for $D_{n}{4000}$ applies to only $\sim 4.4\%$ of the galaxies with $D_{n}4000$ measurements; we removed $\sim 50\%$ of galaxies with initial index values $2.2<D_{n}4000<3$. The average measurement error $\Delta D_{n}4000$ for the final sample is 0.067. In Sect. 2.2, we describe how we used $D_{n}{4000}$ to select quiescent HectoMAP galaxies.